how do i complete the square with a ellipse equation? the equation im having issues with is 49x^2+36y^2+686x-360y+1537=0. please help.
Answer by scott8148(6628)

49(x^2 + 14x) + 36(y^2 - 10y) = -1537

49(x^2 + 14x + 49) + 36(y^2 - 10y + 25) = -1537 + (49 * 49) + (36 * 25)

49(x + 7)^2 + 36(y - 5)^2 = -1537 + (49 * 49) + (36 * 25)

[49(x + 7)^2 + 36(y - 5)^2] / (49 * 36) = [-1537 + (49 * 49) + (36 * 25)] / (49 * 36)